Hi. I am trying to plot 2 fractions and 2 decimals on a number line. They are 1/3, -1/4, -2.5, and 1.0. I have to plot them on a line graph and then put them in order from least to greatest. I am having a lot of trouble with this and hope you can help. Thanks!!!
Answer by KnightOwlTutor(293)

You either have to convert everything to decimal or fractions
1/3, -1/4 , -2 and 5/10=-25/10
The common denomiator for all these is 60
20/60, -15/60,-150/60,60/60
60/60=1
-150/60,-15/60,20/60,60/60
-2.5,-1/4,1/3,1